### v2

# An Efficient Algorithm for Generating Minimal Unique-Cause MC/DC Test cases for Singular Boolean Expressions

Abstract.
Modified Condition/Decision Coverage (MC/DC) is a mandatory structural coverage criterion for ensuring the reliability and safety of critical systems. While its strictest form, Unique-Cause MC/DC, offers the highest assurance, research on its efficient test generation has been lacking. This gap is particularly significant, as an analysis of large-scale avionics systems shows that 99.7% of all conditional decisions are, in fact, Singular Boolean Expressions (SBEs) the ideal structure for applying Unique-Cause MC/DC. This paper proposes ‘Robin’s Rule’, a deterministic algorithm that directly constructs a minimal test set of $N+1$ cases to guarantee 100% Unique-Cause MC/DC for SBEs with $N$ conditions, without generating a full truth table. To validate our approach, we constructed a benchmark by reformulating the TCAS-II specifications into SBEs and verified the results using an industry-standard, certified commercial tool. The results confirm that our method consistently achieves 100% coverage with the theoretical minimum number of tests and is more efficient than the commercial tool. This work provides a practical and provably optimal solution for verifying safety-critical systems, ensuring both rigor and efficiency.
Software Testing, Code Coverage, MC/DC, Singular Boolean Expressions (SBE), Unique-Cause, Test Case Generation
ccs: Software and its engineering Software testing and debugging
1. Introduction
Modern society relies heavily on complex and sophisticated software systems across various domains, including avionics, automotive, medical devices, and industrial control, where a failure can lead to catastrophic human and material losses, making reliability and safety a paramount concern (Martins2020, ), (An2021, ), (Singh2021, ), (Singh2018, ). Within the software development lifecycle, the Verification and Validation (V&V) activities play a critical role in ensuring this reliability, with software testing standing as the primary method for identifying and eliminating potential defects before deployment. However, as software complexity increases, the time and cost required for thorough testing can become prohibitive, often consuming over 50% of a project’s budget and requiring repetitive execution for every code change.
To manage these costs while quantitatively evaluating the effectiveness of testing, the industry widely employs code coverage metrics. Code coverage serves as a measure of test adequacy by indicating the degree to which source code has been exercised by a particular test suite. While criteria like Statement, Branch, and Condition Coverage exist, they often fail to detect intricate logical faults within complex conditional expressions. At the other end of the spectrum, Multiple Condition Coverage (MCC) is the most thorough, requiring $2^{N}$ test cases for $N$ conditions to cover every possible logical combination, but its combinatorial explosion renders it impractical for all but the simplest cases.
Against this backdrop, Modified Condition/Decision Coverage (MC/DC) was established as an effective and practical compromise (Squire2020, ). The core principle of MC/DC is to demonstrate that each individual condition can independently affect the decision’s outcome, regardless of the other conditions. Due to this blend of rigor and efficiency, MC/DC is mandated by numerous international safety standards for software at the highest integrity levels, including DO-178C (avionics) (DO178C, ), ISO 26262 (automotive) (ISO26262, ), IEC 61508 (industrial) (IEC61508, ), and IEC 62304 (medical) (IEC62304, ), EN 50128 (railway) (EN50128, ).
MC/DC is primarily categorized into Unique-Cause and Masking forms (Chilenski2001, ). Unique-Cause MC/DC is the stricter approach that most clearly guarantees a condition’s independence by requiring that only the condition of interest changes its value between a pair of tests. In contrast, Masking MC/DC offers more flexibility by allowing for logical masking like short-circuit evaluation (CAST2001, ). However, this flexibility can lead to subtle issues. For instance, research by Rajan et al.(Rajan2014, ) points out that developers can intentionally alter code structure (e.g., through inlining) to more easily satisfy coverage requirements, a practice sometimes referred to as ‘cheating’. This possibility, arising from the flexibility of Masking MC/DC, underscores the need for research into the more robust Unique-Cause MC/DC for systems that demand the highest level of reliability.
Despite its importance, a significant portion of existing research has focused on Masking MC/DC. This trend was largely driven by the ”solvability problem,” as the strict rules of Unique-Cause MC/DC make it impossible to achieve coverage for many real-world expressions containing coupled conditions (Non-SBEs), where the same condition variable appears more than once. However, this problem is rare in practice. An analysis of large-scale avionics software revealed that 99.7% of all conditional decisions were, in fact, Singular Boolean Expressions (SBEs), with only 70 out of 20,256 being Non-SBEs (Chilenski2001, ). An SBE, an expression where each condition variable appears only once, is the ideal structure for applying Unique-Cause MC/DC to unambiguously prove the independence of each condition. This implies that a highly efficient test generation technique optimized for SBEs would be both practical and highly impactful for the vast majority of real-world safety-critical systems.
Therefore, this paper proposes a new algorithm, ‘Robin’s Rule’, to overcome the limitations of existing test generation methods, such as the exponential complexity of exploring a full truth table ($2^{N}$). Our algorithm directly generates the theoretical minimum of $N+1$ test cases for SBEs, aiming to dramatically reduce verification costs while satisfying the strictest testing standards.
The contributions of this paper are threefold and are detailed as follows:
  1. (1)
    A Deterministic Algorithm for Minimal Test Generation for Unique-Cause MC/DC: We introduce ‘Robin’s Rule’, a deterministic and procedural algorithm that guarantees the theoretical minimum $N+1$ test set for Unique-Cause MC/DC on SBEs. This algorithm overcomes the limitations of existing complex search or constraint-solving methods by analyzing the structural properties of the expression and directly constructing the test cases according to explicit rules.
  2. (2)
    Experimental Proof of Correctness and Optimality: We empirically prove the correctness and optimality of the proposed algorithm. Through extensive experiments on a benchmark dataset based on the real-world safety-critical TCAS-II specification, we show that the generated $N+1$ test cases consistently achieve 100% Unique-Cause MC/DC coverage without exception when validated by a DO-178C certified commercial tool.
  3. (3)
    A Practical Verification Solution for Safety-Critical Systems: We present a practical and cost-effective solution for the verification of safety-critical systems. The results of this study provide a concrete method to dramatically reduce significant testing overhead and costs by reducing the number of test cases to the theoretical minimum, without compromising the strictest reliability standards.
The remainder of this paper is organized as follows. Section 2 discusses background and related work. Section 3 details the proposed algorithm. Section 4 presents the experimental evaluation, and Section 5 discusses the results. Finally, Section 6 concludes the paper.
2. Background and Related Work
2.1. Preliminaries
2.1.1. The Need for and Definition of MC/DC
In software testing, code coverage is a critical metric for test adequacy. While Decision Coverage can miss logical errors within individual conditions, Multiple Condition Coverage (MCC) is often infeasible due to the exponential ($2^{N}$) growth of test cases. MC/DC (Modified Condition/Decision Coverage) was proposed as a practical compromise, requiring that each condition be shown to independently affect the decision’s outcome (Hayhurst2001, ), (Chilenski2001, ). This independence is verified by finding an ‘independence pair’ for each condition: a pair of test cases where only the value of that single condition is toggled, which in turn toggles the final outcome of the overall decision.
This notion of independent effect can be analyzed formally using the Boolean Difference Function, as shown in Equation (1).
(1) $\frac{\delta F(c)}{\delta c_{i}}=F(c_{1},\dots,c_{i},\dots,c_{n})\oplus F(c_{1},\dots,\neg c_{i},\dots,c_{n})$
where $\oplus$ denotes the XOR operation. A result of ‘True’ for this function signifies that toggling the value of $c_{i}$ while all other conditions are held constant will toggle the outcome of the entire function $F(c)$. This function, therefore, serves as the mathematical basis for determining if a condition has an independent influence for a given input vector $c$.
2.1.2. Formal Definitions of Unique-Cause and Masking MC/DC
Using the Boolean Difference Function, the two primary forms of MC/DC can be defined more rigorously (Chilenski2001, ), (CAST2002, ).
Unique-Cause MC/DC
This is the strictest form, requiring that an independence pair, consisting of two test vectors $x$ and $y$ for a condition $c_{i}$, must satisfy all of the following rules (Hayhurst2001, ):
  1. (1)
    $x_{i}\neq y_{i}$
  2. (2)
    $F(x)\neq F(y)$
  3. (3)
    For all $j\neq i$, $x_{j}=y_{j}$
This third rule, ”holding all other conditions fixed,” is the core constraint that ensures the change in condition $c_{i}$ is the ‘unique cause’ for the change in the outcome.
Masking MC/DC
This is a more flexible form that requires an independence pair $x$ and $y$ for a condition $c_{i}$ to satisfy the following (CAST2001, ):
  1. (1)
    $x_{i}\neq y_{i}$
  2. (2)
    $F(x)\neq F(y)$
  3. (3)
    $\frac{\delta F(x)}{\delta x_{i}}=\text{True}\land\frac{\delta F(y)}{\delta y_{i}}=\text{True}$
2.1.3. SBE (Singular Boolean Expressions) and Coupled Conditions
An SBE is defined as a non-degenerate Boolean expression where each condition variable appears only once. In contrast, a Non-SBE such as ‘(A && B) —— (A && C)’ contains Coupled Conditions. Proving the independence of the first instance of ‘A’ while holding the second instance of ‘A’ fixed is logically impossible, making Unique-Cause MC/DC inapplicable. The fact that 99.7% of conditions in large-scale avionics systems are SBEs reinforces the practical importance of this study (Chilenski2001, ).
2.2. Related Work
The automated generation of test cases for MC/DC is a well established research area that has been actively studied over the past decades (Cao2024, ), (Shekhawat2021, ), (Barisal2021, ), (Godboley2017, ), (Kangoye2015, ), (Haque2014, ), (Ghani2009, ), (Awedikian2009, ). Early and theoretical approaches included using the Boolean Difference Function or the Independence Graph model to theoretically prove the feasibility of constructing a minimal $N+1$ test set. Modern research has built upon this foundation, proposing more sophisticated and automated generation techniques.
2.2.1. Formal Model and Constraint Solving-Based Generation
This family of approaches transforms the MC/DC problem into a formal model to be solved by powerful engines. The most active research area involves leveraging SAT/SMT solvers. This technique formalizes all MC/DC rules (e.g., independence pairs) as logical constraints and uses a solver to find a minimal set of test inputs that satisfies them. The works of Kitamura et al.(Kitamura2018, ), Yang et al.(Yang2018, ), and Jaffar et al.(Jaffar2019, ) are representative of this category. While powerful, they often target Masking MC/DC and their internal search process can be a ‘black-box’. Furthermore, model checkers like CBMC have been utilized to explore the program’s state space to find paths satisfying MC/DC (Golla2024, ). Reinforcing this trend, the work by Golla et al.(Golla2025, ) specifically targets the stricter Unique-Cause MC/DC. Their proposed ‘UCM-Gen’ tool generates all possible inequality comparison sequences for a predicate, which are then annotated into the program. The CBMC model checker is subsequently used to formally verify these sequences and compute the UCM Score. This approach highlights a key distinction: while effective, it relies on a powerful, general-purpose verification engine to solve the UCM problem, contrasting with our direct, deterministic algorithm.
Another prominent technique involves transforming Boolean expressions into graph-based data structures. Binary Decision Diagrams (BDDs) represent a Boolean function in a compressed, directed acyclic graph, avoiding the exponential space problem of a full truth table. Test case generation is performed by searching for specific paths on this graph to find combinations that satisfy the independence pairs. However, BDDs are sensitive to variable ordering, and pathfinding often relies on heuristics. The research by Ahishakiye et al. (Ahishakiye2021, ) demonstrated success by combining roBDDs with heuristic pathfinding to generate the $N+1$ minimal set for a significant portion of the TCAS-II benchmark, but did not guarantee minimality for all SBEs.
2.2.2. Search-Based and Heuristic Approaches
These methods frame MC/DC coverage attainment as an optimization problem and employ various search techniques. Search-based heuristics, such as genetic algorithms (Kaur2011, ) or reinforcement learning (Cegin2020, ), are effective in complex search spaces but are often non-deterministic and do not formally guarantee finding the theoretical optimum. Recent work by Sartaj et al. (Sartaj2025, ) has extended this approach into the Model-Based Testing (MBT) domain. They focus on generating MC/DC test data from high-level system constraints written in OCL (Object Constraint Language). Their search-based approach incorporates heuristics like Case-Based Reasoning (CBR) and Range Reduction to improve efficiency. While this presents a practical approach for testing at a higher level of abstraction, it remains fundamentally reliant on search-based heuristics, distinguishing it from our deterministic method.
2.2.3. Code Transformation-Based Approaches
Another interesting line of research involves Code Transformation techniques to facilitate test generation. The ‘SMUP’ technique proposed by Barisal et al.(Barisal2024, ) is a representative example. This method modifies the source code by inserting additional ‘if-else’ blocks with specific patterns around the original conditional statements. The transformed code, while semantically equivalent to the original, is structured to provide new execution paths that guide a concolic testing tool, like jCUTE, to more easily discover MC/DC independence pairs. This work focuses on ‘improving’ MC/DC coverage on existing code, which is a fundamentally different philosophy from our approach of directly generating a minimal set from an unmodified SBE.
2.2.4. Research Trends and Positioning of This Work
There is an important reason why Masking MC/DC became the mainstream approach among these diverse studies. The strict ”hold all other conditions fixed” rule of Unique-Cause MC/DC makes it impossible to achieve coverage for many real-world expressions, especially Non-SBEs that contain coupled conditions (Chilenski2001, ). Due to this ”unsolvable” problem, many researchers and tool developers adopted the more pragmatic Masking MC/DC as a practical alternative with wider applicability (CAST2001, ). However, this focus on pragmatism has, paradoxically, resulted in slowing down research into optimized solutions for SBEs, which constitute 99.7% of actual avionics system code.
Therefore, Robin’s Rule possesses distinct differentiation and originality. Unlike the approaches above that rely on complex external technologies (OCL, SMT, BDDs), heuristics, or code transformation, it provides a deterministic and transparent procedure based on the structural rules of the expression, for the clear scope of SBEs, to directly construct the $N+1$ minimal set from the start without generating a full truth table.
3. The Proposed Algorithm: Robin’s Rule
The core philosophy of Robin’s Rule is to bypass the inefficient process of generating a full truth table ($2^{N}$) and then searching for independence pairs. Our algorithm overcomes this limitation by directly constructing the theoretical minimum of $N+1$ test cases from the start, by analyzing the structural properties and operator relationships of the SBE. This deterministic, multi-phase algorithm is transparent and predictable, without relying on complex external technologies or heuristics. The overall process consists of four main phases as shown in Figure 1.
Figure 1. The four-phase process of Robin’s Rule.
3.1. Phase 1: Preprocessing
The primary goal of this phase is to transform any given SBE into a standardized format to ensure consistent processing. This provides a foundation for the algorithm to apply its rules uniformly. The process is detailed in Algorithm 1.
Algorithm 1 Preprocessing of the SBE
1:Input: An SBE, $S$
2:Output: A rearranged SBE, $R$
3:procedure Pre-Alignment-Process($S$)
4:     if $S$ has a top-level NOT operator then
5:         $R\leftarrow\text{apply\_DeMorgan\_recursively}(S)$triangleright Phase 1.1: Apply De Morgan’s
6:         $R\leftarrow\text{sort\_conditions}(R)$triangleright Proceed to Phase 1.2
7:     else
8:         $R\leftarrow\text{sort\_conditions}(S)$triangleright Phase 1.2: Sort directly
9:     end if
10:     return $R$
11:end procedure
12:
13:function apply_DeMorgan_recursively(expression)
14:     if expression contains NOT then
15:         expression $\leftarrow$ deMorgan(expression) $\triangleright$ Apply to top-level NOT
16:         // Recursively process sub-expressions
17:         expression $\leftarrow$ [apply_DeMorgan_recursively(sub_expr) for sub_expr in expression]
18:     end if
19:     return expression
20:end function
21:
22:function sort_conditions(expression)
23:     // Sort the main conditions by operand count (descending order)
24:     expression $\leftarrow$ sort_by_operand_count(expression)
25:     // Sort sub-conditions within each main condition
26:     expression $\leftarrow$ [sort_conditions(sub_expr) for sub_expr in expression]
27:     return expression
28:end function
  1. (1)
    Normalization of NOT Operations: The algorithm recursively applies De Morgan’s Laws to the SBE. For example, an expression ‘!(A && B)’ is transformed into ‘!A —— !B’. This process ensures that all ‘NOT’ operators apply only to individual conditions (operands), which simplifies subsequent operator analysis.
  2. (2)
    Structural Sorting: The SBE is then sorted. The main logical blocks connected by the top-level operators are arranged in descending order based on their complexity (i.e., the number of operands they contain). This sorting is applied recursively within each sub-expression. This establishes a deterministic processing order, ensuring that more logically complex parts are handled first.
3.2. Phase 2: Temporary Table Generation
Based on the preprocessed SBE, an empty table is created to store the final test cases. For an SBE with $N$ operands, a two-dimensional table of size $(N+1)\times N$ is initialized to generate the theoretical minimum of $N+1$ test cases. Each row in this table corresponds to a single test case, and each column corresponds to a single condition. The specific procedure is described in Algorithm 2. It distinguishes between cases where the number of conditions is two versus more than two. If there are only two conditions, it proceeds directly to test case generation based on the operator (AND/OR).
Algorithm 2 Pseudocode for Generating a Temporary Table
1:Input: A rearranged SBE, $R$
2:Output: A temporary table for storing test cases
3:procedure Temporary_Table_Generation_Process($R$)
4:     Let $N$ be the number of operands in $R$
5:     if $N>2$ then
6:         $rows\leftarrow N+1$
7:         $columns\leftarrow N$
8:         temporary_table $\leftarrow$ generate_table($rows,columns$)
9:         Call_Next_Phases(temporary_table) $\triangleright$ Proceed to Phase 3 & 4
10:     else if $N=2$ then
11:         $rows\leftarrow 3$
12:         $columns\leftarrow 2$
13:         temporary_table $\leftarrow$ generate_table($rows,columns$)
14:         Let $B$ be the operator in $R$
15:         if $B$ is AND then
16:              Generate_Test_Cases(temporary_table, ”AND”)
17:         else if $B$ is OR then
18:              Generate_Test_Cases(temporary_table, ”OR”)
19:         end if
20:     else
21:         return Size_Error
22:     end if
23:end procedure
3.3. Phase 3: Structural Analysis of SBE
In this phase, the structure of the normalized SBE is analyzed to establish the rules and blueprint for deterministically assigning truth values in Phase 4. To achieve this, this section first defines the core terminology and notation used by the algorithm, and then describes the analysis process, which involves decomposing the SBE into structural forms and then generating a relation table.
3.3.1. Terminology and Notation
To systematically analyze any given SBE, Robin’s Rule uses a set of standardized terms and symbols. These are defined in Table 1. This table serves as a glossary for understanding the components of the algorithm described hereafter.
Table 1. Terminology and Notation for SBE Analysis
Symbol Term Description
$Opr_{i}$ Operand An individual condition variable (e.g., A, B, !C).
$i$ Index A sequence number for operands, B-forms, or S-forms, where $i\in\{1,\dots,N\}$.
$o$ Generalized Operator A symbol for a logical AND ($\wedge$) or OR ($\vee$) operator.
$S_{i}$ S-Form The most basic logical unit, consisting of a single operand ($Opr_{i}$).
$B_{i}^{o}$ B-Form A two-operand logical block where two operands are connected by one operator ($B_{i}^{o}=Opr_{i}\circ Opr_{i+1}$).
$Result_{i}$ Result Block An intermediate logical block composed of one or more ‘B’-forms and ‘S’-forms.
$RT$ Relation Table A truth value sequence derived from the SBE’s operator structure.
3.3.2. Decomposition into Structural Forms
First, the algorithm recursively decomposes the input SBE into a combination of the B-forms and S-forms defined in Table 1. For example, an SBE such as ‘(X1 && X2) —— X3’ is recognized as a structure where one ‘B’-form ($B_{1}^{\wedge}=X1\wedge X2$) and one ‘S’-form ($S_{2}=X3$) are combined by an OR operator, represented as $B_{1}^{\wedge}\vee S_{2}$. This step transforms a complex conditional expression into a standardized combination of basic units that the algorithm can process systematically.
3.3.3. Relation Table Generation
Once the structural decomposition is complete, the algorithm analyzes the hierarchy of operators in the decomposed structure to generate a Relation Table (RT). The purpose of this table is to convert the complex, nested operator relationships into a simple, linear sequence of truth values that can be used in Phase 4. For instance, for a structure $B_{1}^{\wedge}\wedge B_{2}^{\vee}$, the algorithm identifies the operator combination (,$\wedge$,$\wedge$,$\vee$ , ∧ , ∨) and, according to predefined rules, converts it into a final truth value sequence, ‘RT(T, T, F)’. This table acts as a clear blueprint that dictates which base patterns to apply and in what order during the test case composition.
3.4. Phase 4: Test Case Generation
In the final phase, the empty table created in Phase 2 is populated with truth values according to the analysis rules established in Phase 3. The core generation mechanism of Robin’s Rule works by analyzing the structure of the SBE and applying the corresponding generation rule. This overall workflow is described in the generalized form in Algorithm 3, where each branch (if-elseif-else) is detailed in the subsequent subsections 3.4.1, 3.4.2, and 3.4.3.
Algorithm 3 Generalized Test Case Generation Procedure based on the Relation Table (RT)
1:Input: $S_{norm}$ (A preprocessed SBE), $RT$ (The Relation Table from Phase 3).
2:Output: $T_{final}$ (The final $(N+1)\times N$ test case table).
3:$SubExprs\leftarrow\textsc{DecomposeSBE}(S_{norm})$
4:$N\leftarrow\textsc{CountOperands}(S_{norm})$
5:$T_{final}\leftarrow\textsc{InitializeTable}(N+1,N)$
[…]
7$\triangleright$ Iterate through SBE blocks, referencing pre-generated RT to expand the table
8:while $SubExprs$ is not empty do
9:     $NextBlock\leftarrow SubExprs.pop()$
10:     $Op\leftarrow\textsc{GetOperatorBetween}(T_{current},NextBlock)$
11$\triangleright$ Table combination and expansion using RT
12:     if $NextBlock$ is a B-Form then
13:         $P_{next}\leftarrow\textsc{GetBasePatterns}(NextBlock)$
14:         $T_{expanded}\leftarrow\textsc{LayoutTables}(T_{current},P_{next})$
[…]
16:     else if $NextBlock$ is an S-Form then
[…]
18:         $\textsc{FillSColumn}(T_{expanded},RT,Op)$triangleright Use RT per Sec. 3.4.2
19:     end if
20$\triangleright$ Handle exceptional cases during composition using RT
21:     if IsExceptionCase($T_{expanded}$then
22:         $T_{corrected}\leftarrow\textsc{HandleException}(T_{expanded},RT)$$\triangleright$ Use RT per Sec. 3.4.3
23:         $T_{current}\leftarrow T_{corrected}$
24:     else
25:         $T_{current}\leftarrow T_{expanded}$
26:     end if
27:end while
28:$T_{final}\leftarrow T_{current}$
29:return $T_{final}$
3.4.1. Test Case Generation for the ‘B o B’ Form
For an SBE with a 4-condition (N=4) structure where two ‘B’-forms are combined, such as $B_{i}^{o}\circ B_{i+1}^{o}$, the algorithm generates a total of $N+1=5$ test cases. The entire process is illustrated in Figure 2 and consists of the following sub-steps:
Figure 2. The overall three-step process for generating N+1 test cases from a ‘B o B’ form.
  1. (1)
    Base Pattern Definition: First, the algorithm utilizes predefined 3-row ‘Base Pattern Tables’ for the fundamental 2-condition ‘B’-form. These patterns are fixed based on operator priority to satisfy Unique-Cause MC/DC. For the ‘AND’ operator, the minimal patterns are (T, T), (T, F), and (F, T). For the ‘OR’ operator, the patterns are (F, F), (F, T), and (T, F). These serve as the essential building blocks for all test case combinations.
  2. (2)
    Pattern Combination using the RT: The algorithm references the RT to combine the base patterns of the two ‘B’-forms. For example, let’s assume the RT derived from the structure $B_{i}^{\wedge}\wedge B_{i+1}^{\vee}$ is (T, T, F). The algorithm maps the first two values (T, T) to $B_{i}^{\wedge}$ and the last two values (T, F) to $B_{i+1}^{\vee}$ to select their respective initial patterns. In this case, (T, T) is selected for $B_{i}^{\wedge}$, and (T, F) is selected for $B_{i+1}^{\vee}$. Afterward, the remaining areas are filled according to the priority test cases defined in (1) Base Pattern Definition.
  3. (3)
    Table Expansion and Completion: The two selected base patterns are used to construct the larger 5x4 test table. To create the final 5-row table, the pattern table for $B_{i+1}^{\vee}$ is placed two rows down relative to the table for $B_{i}^{\wedge}$. The resulting empty regions are then filled deterministically based on the values from the RT(T, T, F). Specifically, the two empty cells in rows 4 and 5 corresponding to $B_{i}^{\wedge}$ are populated with (T, T), the first two values of the RT. Likewise, the two empty cells in rows 4 and 5 for $B_{i+1}^{\vee}$ are populated with (T, F), the last two values of the RT. This systematic process ensures that a complete and minimal set of 5 ($N+1$) test cases, which verifies the independence of all four conditions, is generated without duplication or omission.
3.4.2. Test Case Generation for the ‘B o S’ Form
The process for a 3-condition (N=3) SBE combining a ‘B’-form and a single ‘S’-form differs based on the connecting operator. The core logic is summarized in Figure 3.
  1. (1)
    Case of AND combination (e.g., $B_{i}^{\wedge}\wedge S_{i}$): As illustrated in the examples in Figure 3, this structure may follow a RT of (T, T). The algorithm combines the base pattern table (3x2) for $B_{i}^{\wedge}$ and the pattern (1-column) for $S_{i}$. The core rule is as follows: for the first $N$ test cases, the $S_{i}$ column is populated entirely with T, and only in the final, $N+1$th test case is the value inverted to F to prove the independence of $S_{i}$.
  2. (2)
    Case of OR combination (e.g., $B_{i}^{\wedge}\vee S_{i}$): When combined with OR, the RT changes to (T, F), and the method for filling the $S_{i}$ column is the opposite of the AND case. That is, for the first $N$ test cases, the $S_{i}$ column is populated entirely with F, and only in the final, $N+1$th test case is the value inverted to T to prove the independence of $S_{i}$.
Figure 3. Test case generation process for the ‘B o S’ form.
3.4.3. Exception Handling
For complex SBEs with deeply nested logical blocks, the simple composition rules described previously may be insufficient to guarantee the generation of a minimal $N+1$ test set. Specifically, Exceptional Cases can occur where invalid or duplicate test cases are generated.
  1. (1)
    Defining the Exception Scenario
    An exception can arise when two or more previously generated intermediate Result Blocks, namely $Result_{1}$ and $Result_{2}$, are themselves combined by a logical operator. For example, an exception scenario occurs in a complex structure like the following:
    $(B_{i}^{o}\circ B_{i+1}^{o})\circ(B_{i+2}^{o}\circ S_{i})\dots$
    In this structure, the algorithm broadly recognizes the expression as two intermediate result blocks, such as $Result_{1}=(B_{i}^{o}\circ B_{i+1}^{o})$ and $Result_{2}=(B_{i+2}^{o}\circ S_{i})\dots$ At this final combination point (the ‘Handling Exception’ point), the test patterns generated independently for each ‘Result’ block can conflict, leading to a violation of the $N+1$ rule or causing redundancy.
  2. (2)
    Method for Handling Exceptions: Analysis of the Figure 4 Example
    To handle such exceptions, Robin’s Rule applies a predefined rule as shown in the example in Figure 4 of the source thesis. This figure illustrates the test table generation process for a complex SBE like $({B_{i}}^{\wedge}\vee{B^{\vee}}_{i+1})\wedge({B^{\vee}}_{i+2}\wedge S_{i})$.
    • Here, the RT is derived as (T, F, F, T, F, T). As can be seen at point ① in the figure, a test case duplication occurs when attempting to populate the 5th row using simple composition (i.e., filling with F, T for the ‘B o S’ form).
    • At this moment, the exception handling rule is applied. The arrow at ① first fills the priority pattern (T, F) from the last three values of the RT (T, F, T). Then, at the point indicated by arrow ②, row 6 is filled with the ‘B o S’ form’s RT (F, T), and row 7 is filled with (F, F).
    • Point ③ shows that the ‘B o S’ form’s RT (F, T) is filled. Finally, as shown in Figure 5, the empty region in the bottom-left (rows 6, 7, 8) is populated with the four values from the left side of the RT (T, F, F, T), and the empty region in the top-right (rows 1, 2, 3, 4) is populated with the three values from the right side of the RT (T, F, T) to complete the final test set.
Figure 4. Application of the exception handling rule.
Figure 5. Test case generation
4. Experimental Evaluation
To evaluate the correctness, minimality, and efficiency of the proposed algorithm, ‘Robin’s Rule’, we conducted an experiment designed to answer three key Research Questions (RQs). This chapter details our experimental setup, research questions, and an analysis of the results.
4.1. Experimental Setup
4.1.1. Benchmark Dataset
This experiment is based on the Traffic Collision Avoidance System II (TCAS-II) specification, a well-known and complex real-world benchmark for safety-critical systems (Weyuker1994, ), As the proposed algorithm is optimized for SBEs, we constructed our benchmark dataset by modifying the 15 core logical expressions from the original TCAS-II benchmark into SBE form. This transformation is based on the fact that the vast majority (99.7%) of conditional statements in large-scale avionics software are SBEs (Chilenski2001, ). A comprehensive benchmark suite with a number of conditions ($N$) ranging from 5 to 23 was configured, as detailed in Table 2.
Table 2. TCAS-II SBE specification
No SBE Condition
1 !(a && b) && (c && !d && !e || !f && g && !h || !i && !j && !k) && (l && m && (n || o) && p || q && (r || s) && !t || u && (v || w)) 23
2 a && (!b || !c) && d || e 5
3 a && (!b || !c || d && e && !(!f && g && h && !i || !j && k && l) && !(!m && n && o && p || !q && !r && s)) || t 20
4 (!a && b || c && !d) && !(e && f) && !(g && h) && !(i && j) && ((k && l || m && n) && o && (!p || !q && !r || !s && (!t || !u))) 21
5 (!a && b || c && !d) && !(e && f) && !(g && h) && ((i && j || k && l) && m && (n && o || !p && q)) 17
6 !(a && b) && (!c && d && !e && !f && (g && h || !i && j)) 10
7 a && !b && !c && d && !e && f && (g || !h && (i || j)) && !(k && l || !m && n || o) 15
8 a && !b && !c && !((d && (e || !f && (g || h))) || i && (j || !k && (l || m)) && !n && !o) && !(p && q || !r && s && !t) 20
9 a && !b && !c && (d && (e || !f && (g || h)) && (!i && !j || k ) || !l) && (m && n || !o && p && !q) 17
10 a || b || c || !d && !e && f && g && !h && !i || j && (k || l) && !m 13
11 a && b && (c || d) && e || f && (g || h) && !i || j && (k || l) 12
12 a && (( b || c || d) && e || f && g || h && (i || j || k || l)) || (m || n) && (o || p || q) && r 18
13 (a && b || c && d) && e && (f || (g && (h && i || j && k))) 11
14 (a && b || c && d) && e && (f && g || !h && i) 9
15 !a && b && !c && !d && (e && f || !g && h) 8
4.1.2. Experimental Procedure
The experiment was conducted according to the following systematic procedure:
  1. (1)
    Source Code Generation: For each of the 15 SBEs in the benchmark dataset, a C language source file was automatically generated, containing the respective SBE within an ‘if’ statement.
  2. (2)
    Test Case Generation: Using the implemented Robin’s Rule tool, each C source file was analyzed to generate a set of test cases satisfying Unique-Cause MC/DC, which was then saved in CSV file format. This process was highly efficient, completing in under 5 seconds on average for all SBEs.
  3. (3)
    Coverage Verification: The generated CSV files were imported into the VectorCAST test environment, where the tests were executed and the results were used to measure the Unique-Cause MC/DC coverage.
4.1.3. Validation Tool and Comparison Target
To serve as an oracle for verifying the correctness and coverage of the generated test cases, we selected VectorCAST (ver. 2024). VectorCAST is an industry-standard commercial software that has received Tool Qualification for avionics standards such as DO-178C (Santhanam2007, ), (DO330, ), (Godboley2021a, ). Therefore, its analysis results can serve as a highly reliable baseline. The primary goal of this experiment is to objectively prove the ‘correctness’ of the proposed algorithm, rather than to compete with other tools on performance. The comparison of the number of test cases with VectorCAST is presented as a strong example to demonstrate that the ‘efficiency’ of our algorithm is at a practical level.
4.2. Research Questions (RQs)
This evaluation was structured to answer the following research questions:
  • RQ1
    (Minimality): Can Robin’s Rule consistently generate a minimal test set of size $N+1$ for SBEs with $N$ conditions?
  • RQ2
    (Correctness): Do the $N+1$ test cases generated by our algorithm achieve 100% Unique-Cause MC/DC coverage when verified by a certified tool?
  • RQ3
    (Efficiency): How does the size of the test set generated by Robin’s Rule compare to that generated by an industry-standard commercial tool?
4.3. Case Study: Detailed Analysis of SBE No. 1
This section describes how Robin’s Rule generates a minimal test set, using the benchmark with the most conditions, SBE No. 1 (N=23), as a target.
4.3.1. Phase 1 & 2 Application for SBE No. 1
When the original conditional expression of SBE No. 1 undergoes the Preprocessing phase, it is transformed into the following normalized and sorted structure:
(((((n||o)&&m)&&l)&&p)||(((r||s)&&q)&&!t)||((v||w)
&&u))&&(((c&&!d)&&!e)||%((!f&&g)&&!h)||((!i&&!j)
&&!k))&&(!a||!b)
As there are 23 conditions, a temporary table of size 24 rows by 23 columns is generated.
4.3.2. Test Case Generation Results
In Phase 3, the Relation Table (RT) is derived as RT(F,T,T,T,F,F,T,T,F,F,T,T,T,T,F,T,T,F,T,T,T,F). Then, in Phase 4, Robin’s Rule uses this RT to generate the final set of 24 ($N+1$) test cases for SBE No. 1, which is the theoretical minimum, as shown in Table 3. Each row represents a single test case, and each column represents the truth value (1=True, 0=False) for each condition from ‘a’ to ‘w’.
Table 3. The 24 Test Cases Generated by Robin’s Rule for SBE No. 1 (N=23).
Test Case a b c d e f g h i j k l m n o p q r s t u v w
01 0 1 1 1 1 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
02 0 0 1 1 1 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
03 1 0 1 1 1 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
04 0 1 0 1 1 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
05 0 1 1 0 1 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
06 0 1 1 1 0 0 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
07 0 1 1 1 0 0 1 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
08 0 1 1 1 0 1 0 1 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
09 0 1 1 1 0 0 1 0 0 0 0 1 1 0 0 1 1 0 1 0 0 0 1
10 0 1 1 1 0 0 1 1 1 0 0 1 1 0 0 1 1 0 1 0 0 0 1
11 0 1 1 1 0 0 1 1 1 0 1 1 1 0 0 1 1 0 1 0 0 0 1
12 0 1 1 1 0 0 1 1 1 1 0 1 1 0 0 1 1 0 1 0 0 0 1
13 0 1 1 1 0 0 1 1 1 0 1 0 1 0 0 1 1 0 1 0 0 0 1
14 0 1 1 1 0 0 1 1 1 0 1 1 1 1 0 1 1 0 1 0 0 0 1
15 0 1 1 1 0 0 1 1 1 0 1 1 0 0 0 1 1 0 1 0 0 0 1
16 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 1 1 0 1 0 0 0 1
17 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 0 1 0 0 0 1
18 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 0 0 1 0 0 0 1
19 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 1 0 0 0 1
20 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 0 0 0 0 1
21 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 0 1 0 0 1
22 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 0 0 1 0 1
23 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 0 0 0 1 1
24 0 1 1 1 0 0 1 1 1 0 1 1 1 0 1 0 1 1 0 0 0 1 0
4.4. Threats to Validity
To ensure the academic rigor of our study, we identify and discuss the potential threats to the validity of our findings.
(Construct Validity)
This threat relates to whether we are truly measuring what we claim to be measuring. Our study relies on VectorCAST, an industry-standard commercial tool, as the oracle for MC/DC coverage. While this tool is certified for DO-178C, which strengthens our confidence, there remains a theoretical possibility that its specific implementation or interpretation of the Unique-Cause criterion could influence the results. To mitigate this, we chose a tool that is widely accepted and trusted in the safety-critical domain.
(Internal Validity)
This threat concerns potential flaws within our own experimental setup. There could be an error in our implementation of the Robin’s Rule algorithm. We mitigated this threat by cross-validating our results: for every SBE, the test set generated by our tool was independently verified by an external, certified tool (VectorCAST) to achieve 100% coverage, which provides strong evidence for the correctness of our implementation.
(External Validity)
This threat relates to the generalizability of our results. First, our benchmark, while based on the TCAS-II specification, consists of SBEs that were manually modified from the original expressions. It is possible that this transformation simplified some of the original logic. Second, our study is intentionally focused on SBEs. While our introduction provides a strong rationale for this focus (99.7% prevalence), the results cannot be directly generalized to Non-SBEs containing coupled conditions. Extending the algorithm to handle Non-SBEs remains a crucial area for future work.
4.5. Overall Result and Analysys
This section analyzes the experimental results for each research question (RQ) in sequence.
Answering RQ1: Generation of Minimal N+1 Test Sets
To answer RQ1, we analyzed the number of test cases generated by Robin’s Rule. For every SBE specification with $N$ conditions, our algorithm generated exactly $N+1$ test cases. This result was consistent across the entire benchmark, from N=5 to N=23, empirically confirming that Robin’s Rule successfully generates a test set of the theoretical minimum size for Unique-Cause MC/DC.
Answering RQ2: Achievement of 100% Coverage
To answer RQ2, each of the $N+1$ test sets was validated using VectorCAST. The results show that 100% Unique-Cause MC/DC coverage was achieved in all 15 cases. This demonstrates the correctness and effectiveness of the proposed deterministic generation logic.
Answering RQ3: Comparison with Commercial Tool
To address RQ3, we compared the number of test cases generated by our algorithm to the results from VectorCAST’s own test case generator. Robin’s Rule consistently generated a smaller test set in every case. The reduction in test cases ranged from one to six, with an average reduction of 2.87 test cases per SBE. This clearly demonstrates the practical efficiency and optimality achieved by the proposed deterministic approach compared to existing industry-grade tools.
In summary, our experiment provides strong evidence that Robin’s Rule is a correct, effective, and efficient method for generating a minimal test set for Unique-Cause MC/DC for SBEs. The detailed data for these results can be found in Table 4.
Table 4. Test Generation Results: Robin’s Rule vs. VectorCAST.
SBE No. of VectorCAST Robin’s Rule
No. Cond. ($N$) Cases Coverage Cases Coverage
1 23 27 100% 24 100%
2 5 8 100% 6 100%
3 20 26 100% 21 100%
4 21 28 100% 22 100%
5 17 22 100% 18 100%
6 10 13 100% 11 100%
7 15 17 100% 16 100%
8 20 24 100% 21 100%
9 17 20 100% 18 100%
10 13 16 100% 14 100%
11 12 15 100% 13 100%
12 18 22 100% 19 100%
13 11 15 100% 12 100%
14 9 12 100% 10 100%
15 8 10 100% 9 100%
5. Discussion
The results have three main implications: optimality and correctness, efficiency in computational complexity, and deterministic reliability. Robin’s Rule avoids the exponential cost of truth-table or exhaustive search methods by directly constructing the minimal set. Its deterministic nature is a key advantage for tool qualification where predictability and reproducibility are paramount. The primary threat to validity is the study’s focus on SBEs. While SBEs are highly prevalent (Chilenski2001, ), extending this work to Non-SBEs is necessary. Another limitation is the comparison against a single commercial tool. VectorCAST was chosen as a certified oracle to primarily validate correctness, and a broader comparative study against other tools remains an area for future work.
6. Conclusion and Future Work
This paper presented Robin’s Rule, a deterministic algorithm that directly constructs a minimal $N+1$ test set for SBEs to satisfy Unique-Cause MC/DC. Experiments on a TCAS-II based benchmark, validated with a certified tool, proved that our method consistently achieves 100% coverage with the theoretical minimum number of test cases. This work offers a practical and optimal solution for developers in safety-critical domains, enabling rigorous software verification at a reduced cost. Future work will focus on extending the algorithm to handle Non-SBEs (coupled conditions) and adapting it for the Masking MC/DC environment.